### v1

@@ Abstract
We consider routing in reconfigurable networks, which is also known as coflow scheduling in the literature. The algorithmic literature generally (perhaps implicitly) assumes that the amount of data to be transferred is large. Thus the standard way to model a collection of requested data transfers is by an integer demand matrix $D$, where the entry in row $i$ and column $j$ of $D$ is an integer representing the amount of information that the application wants to send from machine/node $i$ to machine/node $j$. A feasible coflow schedule is then a sequence of matchings, which represent the sequence of data transfers that covers $D$. In this work, we investigate coflow scheduling when the size of some of the requested data transfers may be small relative to the amount of data that can be transferred in one round. fractional matchings and/or that employ indirect routing, and compare the relative utility of these options. We design algorithms that perform much better for small demands than the algorithms in the literature that were designed for large data transfers.
@@ 1 Introduction
We consider routing in reconfigurable networks—also known as coflow scheduling—whose topologies dynamically adapt to the structure of the communication traffic [CS12a, AS25a]. There is currently no consensus in the community on what the best reconfigurable network design is, and the community lacks models and metrics to rigorously study and compare different designs [AS25a]. Our high level goal in this work is to fill some of this gap.
@@ 1.1 Background and Motivation
The algorithmic literature in this area generally models this informal goal by the formal constraint that at all times the pairs of nodes communicating form a matching $M$, which in this context is a collection $M=\{(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{n},t_{n})\}$ of source-destination pairs of nodes with distinct sources and distinct sinks (so $i\neq j$ implies $s_{i}\neq s_{j}$ and $t_{i}\neq t_{j}$). A coflow schedule is then a sequence of matchings $M_{1},\ldots,M_{T}$, where matching $M_{t}$ is passed to the networking layer at time $t$, and during the time interval $[t,t+1]$ the network layer transmit a unit of data from $s_{i}$ to $t_{i}$ for each $(s_{i},t_{i})\in M_{t}$.
@@
In the algorithmic literature on coflow scheduling [CS15a, CZS14a, Ahm+20a, Cho+19a, Im+19a, RS25a, JKR17a, QSZ15a, DM20a] the input passed to the coflow layer is assumed to consist of an integer demand matrix $D$, where $D_{ij}$, the entry in row $i$ and column $j$ of $D$, is an integer representing the amount of information that the application wants to send from machine/node $i$ to machine/node $j$. Then generally the objective is to optimize some standard quality-of-service metric, like minimizing makespan, which is the time when the last unit of data is transmitted, or minimizing the average completion (or delay) time, which is the average time that it takes a unit of data to reach its destination. So, in some sense, this is modeling scheduling coflows for a single application. (Some of the coflow scheduling literature also considers generalizations of this basic setting, see Section˜1.3.).
@@
The algorithmic literature generally (perhaps implicitly) assumes that the number of requests for data transfers, and the size of these transfers, is large, as this was the situation that was the original motivation for the proposal of a coflow layer [CS12a]. This is reflected in the modeling assumption that the demands $D_{ij}$ are integer. The goal of our research was to examine coflow scheduling when one removes the assumption in the algorithmic coflow literature that the amount of data requested to be transferred in each requested data transfer is large. This assumption is not appropriate in settings where many of the data transfers are small, and it is certainly at least conceivable that such a situation could sometimes arise in data-parallel applications. We model this by allowing the demands $D_{ij}$ to be arbitrary nonnegative rational numbers, and in particular a demand $D_{ij}$ may be less than 1.
@@
To understand the additional issues that arise if some of the requested data transfers are small, consider the demand matrix $D$ where each entry $D_{ij}=\frac{1}{2n^{3/2}}$, where $n$ is the number of nodes used by the application. Then these data transfers can be supported by routing up to a unit of data from node $i$ to node $j$ at the time $t$ where $t=(j-i)\mod n$. This results in a makespan of n1n-1 (and average completion/delay time of (n1)/2(n-1)/2). And it might seem that obtaining sublinear makespan is impossible, as $D$ cannot be expressed as a linear combination of less than n1n-1 matchings.
@@
However, makespan $O(\sqrt{n})$ can be achieved if the coflow manager can indirectly route data (which is commonly assumed to be allowable in the reconfigurable network routing literature [Shr+19a, Bal+20a, Ami+24a, Gho+17a, Mel+24a, AAS23a]. As an example of indirect routing, a unit of data with source $s_{i}$ and destination $t_{i}$ could be routed using two matchings $M_{1}$ and $M_{2}$. So $M_{1}$ could send this data from source node $s_{i}$ to an intermediate node $k$, and $M_{2}$ could route this data from node $k$ to destination node $t_{i}$. One can achieve $O(\sqrt{n})$ makespan (for the instance where each $D_{ij}=\frac{1}{2n^{3/2}}$) in the following way. The $n$ nodes are conceptually assigned to nodes in an $\sqrt{n}$ by $\sqrt{n}$ grid square $G$. Then the data transfers are divided into two phases, and each phase is divided into $\sqrt{n}-1$ subphases. In subphase $k$ of Phase 1, the node in row $i$ column $j$ of $G$ routes to the node in row $(i+k)\mod\sqrt{n}$ column $j$ all its data whose final destination is in row $(i+k)\mod\sqrt{n}$. Thus after the first phase, the current location of each unit of data is in the same row as its final destination. Then subphase $k$ of Phase 2, the node in row $i$ column $j$ of $G$ routes to the node in row $i$ column $(j+k)\mod\sqrt{n}$ all its data whose final destination is that node. See Figure˜1 for an illustration of this algorithm when $n=9$.
@@
Figure 1: The data transmissions for nodes in positions $(0,0)$ and $(1,1)$ for the 3 by 3 grid formed from $n=9$ nodes. The matchings $M_{1}$ and $M_{2}$ belong to Phase 1, and the matchings $M_{3}$ and $M_{4}$ belong to Phase 2.
@@
Here we consider eight variations of coflow scheduling with rational demand matrices, where eight arises from there being two options to each of the following three characteristics:
@@
Makespan vs. Average Completion Times Objective.
@@
One conceptual “solution” for coflow scheduling of small data transfers is to just change the assumptions about the networking layer to assuming that the network layer can route a fractional matching in each unit of time. In this context, a fractional matching is given by a collection $\{(s_{1},t_{1},p_{1}),(s_{2},t_{2},p_{2}),\ldots,(s_{n},t_{n},p_{n})\}$ of triples $(s_{i},t_{i},p_{i})$ indicating that $p_{i}$ fractional units of data are routed from $s_{i}$ to $t_{i}$, such that for all nodes $v$, $\sum_{i:s_{i}=v}p_{i}\leq 1$ and $\sum_{i:t_{i}=v}p_{i}\leq 1$. The relaxation to fractional matchings is at least defensible in settings where the motivation for restricting the transfers to be matchings arises from throughput constraints from the networking layer, as a fractional matching transfers the same amount of data per unit time as an integer matching. In contrast, this relaxation is probably not defensible in settings where the network is optical and the matching is being implemented by the rotation of mirrors [AS25a, Liu+23a, Mel+24a]. Thus we consider both allowing fractional matchings, and restricting matchings to be integral.
So the standard algorithmic research approach would be to, for each of these eight variations, find the best approximation ratio that is achievable by a polynomial-time algorithm. Thus let us start by noting what can be immediately deduced about the best achievable approximation ratio from known results. These observations are summarized in Table˜1.
@@
makespan 1 1 1 $O(\log n)$
average completion 1 $O(\log n)$ $\sqrt{2}$ [GM09a] $O(\log n)$
Table 1: Overview of known approximation guarantees.
For direct integral routing for the makespan objective, optimal schedules can be computed using an optimal algorithm for edge coloring a bipartite graph [Sch+03a]. For fractional routing with the makespan objective, the optimal makespan is the same for both direct and indirect routing because the fractional edge coloring number for a bipartite graph is the same as the maximum coloring number of the two underlying bipartite graphs [Sch+03a]. Finally it is straightforward to model fractional direct routing to optimize average completion time with a polynomial-size linear program. A $O(\log n)$ approximation is achievable for the rest of the variants by modeling the problem as a linear program, where the edges in each time step are over-capacitated by $O(\log n)$. It can be established that this over-capacitation is sufficient via the probabilistic method. For direct integral routing for average completion time, rounding the demands up to the next integer essentially does not change the problem, one can use the $\sqrt{2}$-approximation for integer demands is also $\sqrt{2}$-approximate for rational demands [GM09a].
In this paper, we focus on the variants where a $O(1)$ approximation does not (to the best of our knowledge) follow directly from known results, which all involve indirect routing.
@@ 1.2 Summary of Our Results
We present two main results. Theorem 1 presents the first $O(1)$-approximation algorithm for indirect fractional routing for the average completion time objective.
@@
There is a polynomial-time indirect routing fractional matching algorithm $\mathrm{ALG}$} that outputs, for any valid instance, a schedule with an average completion time that is at most 16 times the optimal average completion time. The algorithm only uses direct routing schedules, so the approximation bound is valid both for the direct and indirect routing cases using fractional matchings.
@@
Algorithm $\mathrm{ALG}$} from Theorem˜2 attains a $O(\log n)$-approximation ratio for both makespan and average completion time, for all $B$.
In some sense, the challenge of obtaining a $O(1)$-approximation appears to be to identify instances that admit schedules with less than worst-case objective value.
If $B\leq 2$, we can use Valiant’s hypercube design [Val82a] to achieve $O(\log n)$ makespan by using routing paths with $O(\log n)$ intermediate nodes. In a nutshell, this approach tries to maximize the usage of intermediate nodes, because there is very little demand to route. Accordingly, the $\Omega(\log n)$ makespan lower bound relies purely on topological properties, and does not take demands into account. Alternatively, if $B\geq n$, we can use a single round-robin design and route flow directly from sources to destinations in maximum matchings to achieve $O(B)$ makespan—there is no need for intermediate nodes because there are no small demands that can be “combined.” The $\Omega(B)$ lower bound relies purely on the maximum demand requested per source or destination node, without taking the network topology into account.
@@ 1.3 Related Work
Coflow scheduling was popularized by [CS12a], and followed by heuristics for online and offline settings [CZS14a, CS15a]. A recent overview of reconfigurable networking can be found in [AS25a].
@@
There is some literature on direct integral routings to optimize weighted completion time. Marx [Mar09a] showed that this problem is APX-hard. The natural greedy algorithm is a 2-approximation [Bar+98a, Bar+00a]. [GM09a] give an improved $\sqrt{2}$-approximation, and [HKS11a] show a $1.83$-approximation for a more general setting. Other works in this area consider non-preemptive variants, where large demands must be scheduled consecutively in time [HKS01a, Kim05a, Gan+08a]. The fractional preemptive variants of these problems are also captured by the general polytope scheduling problem, as any feasible instantaneous allocation belongs to the matching polytope [IKM18a, JLM25a].
@@
There is also some literature on the situation where there are several different applications, each with their own demand matrix, and the objective is to minimize the sum of the makespans over the different applications. This literature considers direct integral routing for the average completion time objective. The problem is APX-hard [Mar09a], and an even stronger inapproximability of better-than-2 follows from a reduction to the concurrent open shop problem [QS02a], which is hard under the Unique Games Conjecture [BK09a]. The first constant-factor approximation was shown by [QSZ15a], and after several improvements [KP16a, SG17a, Ahm+20a], the current best approximation ratio is $3.415$ [RS25a]. The special case where each coflow corresponds to the set of incident edges of a vertex is called data migration and also has been studied extensively [Kim05a, KKW04a, Gan+08a, GM09a, Mes10a]. Other notable results are online algorithms [Khu+19a, BNV20a, DM20a], a matroid generalization [Im+19a], and generalization for general graphs [Cho+19a, JKR17a]. In particular some elements of our dual fitting analysis derive from the dual fitting analysis in [DM20a].
@@
The second main area of related work is in Oblivious Reconfigurable Networks (ORNs). The theoretical problem was formalized recently [Ami+22a], modeling the constraints of various proposed data center architectures using fast circuit switch technology [Gho+17a, Mel+24a, Shr+19a, AAS23a, Ami+24a]. The ORN model allows for multi-hop indirect routing, its key connection to our work. However, it also focuses on the streaming setting, where traffic demands arrive over time and must be continuously fulfilled. Additionally, they focus on obliviousness, and frame their results as a guarantee on the minimum amount of traffic which can be fulfilled, rather than as a guaranteed approximation ratio. One contribution of the ORN area is a generalized hypercube-style design with a tunable hop parameter [Ami+22a, Wil+23a, Ami+24a]. Our work in Section˜4 uses similar ideas, but applied to a different use case. Randomized oblivious connection schedules with some adaptive routing have also been considered in theory [Wil+24a], focusing on the same tradeoff guarantees as in the oblivous literature. It is still unknown what can be done with a fully adaptive design.

### v2

@@ Abstract
We consider routing in reconfigurable networks, which is also known as coflow scheduling in the literature. The algorithmic literature generally (perhaps implicitly) assumes that the amount of data to be transferred is large. Thus the standard way to model a collection of requested data transfers is by an integer demand matrix $D$, where the entry in row $i$ and column $j$ of $D$ is an integer representing the amount of information that the application wants to send from machine/node $i$ to machine/node $j$. A feasible coflow schedule is then a sequence of matchings, which represent the sequence of data transfers that covers $D$. In this work, we investigate coflow scheduling when the size of some of the requested data transfers may be small relative to the amount of data that can be transferred in one round. In particular, we investigate algorithms that employ fractional matchings and/or that employ indirect routing, and compare the relative utility of these options. We design algorithms that perform much better for small demands than the algorithms in the literature that were designed for large data transfers.
@@ 1 Introduction
We consider routing in reconfigurable networks, also known as coflow scheduling, whose topologies dynamically adapt to the structure of the communication traffic [12, 5]. There is currently no consensus in the community on what the best reconfigurable network design is, and the community lacks models and metrics to rigorously study and compare different designs [5]. Our high level goal in this work is to fill some of this gap.
@@ 1.1 Background and Motivation
The algorithmic literature in this area generally models this informal goal by the formal constraint that at all times the pairs of nodes communicating form a matching $M$, which in this context is a set $M=\{(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{n},t_{n})\}$ of source-destination pairs of nodes with distinct sources and distinct sinks (so $i\neq j$ implies $s_{i}\neq s_{j}$ and $t_{i}\neq t_{j}$). A coflow schedule is then a sequence of matchings $M_{1},\ldots,M_{T}$, where matching $M_{t}$ is passed to the networking layer at time $t$, and during the time interval $[t,t+1]$ the network layer transmits a unit of data from $s_{i}$ to $t_{i}$ for each $(s_{i},t_{i})\in M_{t}$.
@@
The input passed to the coflow layer is assumed to consist of an integer demand matrix $D$, where $D_{ij}$, the entry in row $i$ and column $j$ of $D$, is an integer representing the amount of information that the application wants to send from machine/node $i$ to machine/node $j$ (see, e.g., [13, 14, 2, 11, 23, 37, 25, 35, 15]). Then the objective is to optimize some standard quality-of-service metric, like minimizing makespan, which is the time when the last unit of data is transmitted, or minimizing the average completion (or delay) time, which is the average time that it takes a unit of data to reach its destination. Some of the coflow scheduling literature also considers generalizations of this basic setting, see Section˜1.3.
@@
Originally, the proposal of a coflow layer was to handle large numbers of data transfer requests, whose sizes were also large [12], and so this has been the, perhaps implicit, assumption in most of the algorithmic literature. This is reflected in the modeling assumption that the demands $D_{ij}$ are integer. Our goal in this work was to examine coflow scheduling when the amount of data transfer requests may be small, as one could conceive in certain data-parallel applications. We model this by allowing the demands $D_{ij}$ to be arbitrary nonnegative rational numbers; in particular demands may be less than 1.
@@
To understand the additional issues that arise if some of the requested data transfers are small, consider the demand matrix $D$ where each entry $D_{ij}=\frac{1}{2n^{3/2}}$, where $n$ is the number of nodes used by the application. Then these data transfers can be supported by routing up to a unit of data from node $i$ to node $j$ at the time $t$ where $t=(j-i)\mod n$. This results in a makespan of n1n-1 (and average completion time of (n1)/2(n-1)/2), and it might seem that obtaining sublinear makespan is impossible, as $D$ cannot be expressed as a linear combination of less than n1n-1 matchings.
@@
However, a makespan of $O(\sqrt{n})$ can be achieved if the coflow scheduler can indirectly route data, as it is commonly assumed in the reconfigurable networks literature [40, 6, 3, 18, 33, 1]. As an example of indirect routing, a unit of data with source $s_{i}$ and destination $t_{i}$ could be routed using two matchings $M_{1}$ and $M_{2}$: $M_{1}$ could send the data from source node $s_{i}$ to an intermediate node $k$, and $M_{2}$ could route this data from node $k$ to destination node $t_{i}$. One can achieve a makespan of $O(\sqrt{n})$ (for the instance where each $D_{ij}=\frac{1}{2n^{3/2}}$) in the following way. The $n$ nodes are conceptually assigned to nodes in an $\sqrt{n}\times\sqrt{n}$ square grid $G$. Then the data transfers are divided into two phases, and each phase is divided into $\sqrt{n}-1$ subphases. In subphase $k$ of Phase 1, the node in row $i$ and column $j$ of $G$ routes to the node in row $(i+k)\mod\sqrt{n}]$ and column $j$ all of its data whose final destination is in row $(i+k)\mod\sqrt{n}]$. Thus after the first phase, the current location of each unit of data is in the same row as its final destination. Then subphase $k$ of Phase 2, the node in row $i$ and column $j$ of $G$ routes to the node in row $i$ and column $(j+k)\mod\sqrt{n}]$ all of its data whose final destination is that node. See Figure˜1 for an illustration of this algorithm when $n=9$.
@@
Figure 1: The data transmissions for nodes $(0,0)$ and $(1,1)$ for the $3\times 3$ grid formed from $n=9$ nodes. The matchings $M_{1}$ and $M_{2}$ belong to Phase 1, and the matchings $M_{3}$ and $M_{4}$ belong to Phase 2.
@@
Here we consider eight variations of coflow scheduling with rational demand matrices, where “eight” comes from there being two options to each of the following three characteristics:
@@
Makespan vs. Average Completion Time Objective.
@@
One conceptual “solution” for coflow scheduling of small data transfers is to allow the network layer to route a fractional matching in each unit of time. In this context, a fractional matching is given by a collection $\{(s_{1},t_{1},p_{1}),(s_{2},t_{2},p_{2}),\ldots,(s_{n},t_{n},p_{n})\}$ of triples $(s_{i},t_{i},p_{i})$ indicating that $p_{i}$ fractional units of data are routed from $s_{i}$ to $t_{i}$, such that for all nodes $v$, $\sum_{i:s_{i}=v}p_{i}\leq 1$ and $\sum_{i:t_{i}=v}p_{i}\leq 1$. Note that a fractional matching transfers the same amount of data per unit time as an integer matching, so it does not violate the unit throughput constraints at the network layer and thus it would be feasible in several networking scenarios. In contrast, such a relaxation is probably not feasible in settings where the network is optical and the matching is being implemented by the rotation of mirrors [5, 31, 33]. Thus we consider both allowing fractional matchings and restricting matchings to be integral.
For each of these eight variations, the goal is to find the best approximation ratio that is achievable by a polynomial-time algorithm. Table˜1 summarizes the best-known approximation ratios for all the variations of coflow scheduling.
Table 1: Overview of known approximation guarantees. We have provided references to some of the results, as appropriate; for the others, the corresponding approximation bounds follow directly from well-known standard approximation techniques.
@@
makespan [38] [38] 1 $O(\log n)$
average completion 1 $O(\log n)$ $\sqrt{2}$ [17] $O(\log n)$
For direct integral routing for the makespan objective, optimal schedules can be computed using an optimal algorithm for edge coloring a bipartite graph [38]. For fractional routing with the makespan objective, the optimal makespan is the same for both direct and indirect routing because the fractional edge coloring number for a bipartite graph is the same as the maximum coloring number of the two underlying bipartite graphs [38]. Finally it is straightforward to model fractional direct routing to optimize average completion time with a polynomial-size linear program. An $O(\log n)$-approximation is achievable for the rest of the variants by modeling the problem as a linear program, where the edges in each time step are over-capacitated by $O(\log n)$ [19]. It can be established that this over-capacitation is sufficient via the probabilistic method. For direct integral routing for average completion time, rounding the demands up to the next integer essentially does not change the problem, one can use the $\sqrt{2}$-approximation for integer demands is also $\sqrt{2}$-approximate for rational demands [17].
In this paper, we focus on the variants where a $O(1)$-approximation does not (to the best of our knowledge) follow directly from known results, which all involve indirect routing.
@@ 1.2 Summary of Our Results
We present two main results, stated in Theorems 1 and 2. Theorem 1 presents the first $O(1)$-approximation algorithm for indirect fractional routing for the average completion time objective:
@@
There is a polynomial-time algorithm $\mathrm{ALG}$} for indirect routing with fractional matchings that outputs, for any valid instance, a schedule whose average completion time is at most 16 times the optimal average completion time. The algorithm only uses direct routing schedules, so the approximation bound is valid for both the direct and indirect routing settings with fractional matchings.
@@
Algorithm $\mathrm{ALG}$} from Theorem˜2 attains an $O(\log n)$-approximation ratio for both makespan and average completion times, for all $B$.
In some sense, the challenge of obtaining an $O(1)$-approximation appears to be to identify instances that admit schedules with less than worst-case objective value.
If $B\leq 2$, we can use Valiant’s hypercube design [42] to achieve $O(\log n)$ makespan by using routing paths with $O(\log n)$ intermediate nodes. In a nutshell, this approach tries to maximize the usage of intermediate nodes, because there is very little demand to route. Accordingly, the $\Omega(\log n)$ makespan lower bound relies purely on topological properties, and does not take demands into account. Alternatively, if $B\geq n$, we can use a single round-robin design and route flow directly from sources to destinations in maximum matchings to achieve $O(B)$ makespan—there is no need for intermediate nodes because there are no small demands that can be “combined.” The $\Omega(B)$ lower bound relies purely on the maximum demand requested per source or destination node, without taking the network topology into account.
@@ 1.3 Related Work
Coflow scheduling was popularized by Chowdhury and Stoica [12], and followed by heuristics for online and offline settings [14, 13]. A recent overview of reconfigurable networking can be found in [5].
@@
There is some literature on direct integral routings to optimize weighted completion time. Marx [32] showed that this problem is APX-hard. The natural greedy algorithm is a 2-approximation [8, 9]. Gandhi and Mestre [17] give an improved $\sqrt{2}$-approximation, and Halldórsson, Kim, and Sviridenko [21] show a $1.83$-approximation for a more general setting. Other works in this area consider non-preemptive variants, where large demands must be scheduled consecutively in time [20, 29, 16]. The fractional preemptive variants of these problems are also captured by the general polytope scheduling problem, as any feasible instantaneous allocation belongs to the matching polytope [22, 24].
@@
There is also some literature on the situation where there are several different applications, each with their own demand matrix, and the objective is to minimize the sum of the makespans over the different applications. This literature considers direct integral routing for the average completion time objective. The problem is APX-hard [32], and an even stronger hardness-of-approximation result ruling out any factor better than $2$ follows from a reduction to the concurrent open shop problem [36], which is hard under the Unique Games Conjecture [7]. The first constant-factor approximation was shown by Qiu, Stein, and Zhong [35], and after several improvements [28, 39, 2], the current best approximation ratio is $3.415$ [37]. The special case where each coflow corresponds to the set of incident edges of a vertex is called data migration and also has been studied extensively [29, 26, 16, 17, 34]. Other notable results are online algorithms [27, 10, 15, 30], a matroid generalization [23], and generalization for general graphs [11, 25]. In particular some elements of our dual fitting analysis derive from the dual fitting analysis in [15].
@@
The second main area of related work is in Oblivious Reconfigurable Networks (ORNs). The theoretical problem was formalized recently [4], modeling the constraints of various proposed data center architectures using fast circuit switch technology [18, 33, 40, 1, 3]. The ORN model allows for multi-hop indirect routing, its key connection to our work. However, it also focuses on the streaming setting, where traffic demands arrive over time and must be continuously fulfilled. Additionally, they focus on obliviousness, and frame their results as a guarantee on the minimum amount of traffic which can be fulfilled, rather than as a guaranteed approximation ratio. One contribution of the ORN area is a generalized hypercube-style design with a tunable hop parameter [4, 44, 3]. Our work in Section˜4 uses similar ideas, but applied to a different use case. Randomized oblivious connection schedules with some adaptive routing have also been considered in theory [43], focusing on the same tradeoff guarantees as in the oblivious literature. It is still unknown what can be done with a fully adaptive design.